OpenStudy (blm1):

I am just trying to figure out how many targets they are making. How do I figure this out? The seventh-grade class is building target areas for a PE activity. The bases for the game will be in a circular shape. The diameter of each circle is 3 feet. Approximately, how many square feet of the turf need to be painted for a base circle? Use 3.14 for π and round your answer to the nearest tenth. 7.1 square feet 9.4 square feet 18.8 square feet 28.3 square feet

OpenStudy (phi):

It does not matter how many targets they are making (to answer their question) the question is what is the area of *one* circle I would use Area= pi*r*r with r= 3/2 (diameter divided by 2)

OpenStudy (blm1):

Not to be rude but I am still confused. I don't understand what I am supposed to be doing.